### v1

# TopoBenchmarkX: A Framework for Benchmarking Topological Deep Learning

Abstract
This work introduces TopoBenchmarkX, a modular open-source library designed to standardize benchmarking and accelerate research in Topological Deep Learning (TDL). TopoBenchmarkX maps the TDL pipeline into a sequence of independent and modular components for data loading and processing, as well as model training, optimization, and evaluation. This modular organization provides flexibility for modifications and facilitates the adaptation and optimization of various TDL pipelines. A key feature of TopoBenchmarkX  is that it allows for the transformation and lifting between topological domains. This enables, for example, to obtain richer data representations and more fine-grained analyses by mapping the topology and features of a graph to higher-order topological domains such as simplicial and cell complexes. The range of applicability of TopoBenchmarkX is demonstrated by benchmarking several TDL architectures for various tasks and datasets.
1 Introduction
In geometric deep learning (GDL; Bronstein et al., 2021), graph neural networks (GNNs; Zhou et al., 2020) have demonstrated impressive capabilities in processing relational data represented as graphs. However, graphs only encode pairwise structures, which can be limiting. For example, social interactions may involve multiple individuals, and electrostatic interactions in proteins may involve multiple atoms. Topological deep learning (TDL; Papamarkou et al., 2024; Bodnar, 2023; Hajij et al., 2023; Papillon et al., 2023b) offers a framework for modeling complex systems characterized by such multi-way (higher-order) relations among components. Topological neural networks (TNNs; Feng et al., 2019; Bunch et al., 2020; Hajij et al., 2020; Bodnar et al., 2021a; Ebli et al., 2020; Schaub et al., 2021; Bodnar et al., 2021b; Chien et al., 2021), which are part of TDL, have found applications in numerous fields that involve higher-order relational data such as social networks (Knoke and Yang, 2019) and protein biology (Jha et al., 2022). TNNs achieve state-of-the-art performance in various machine learning tasks (Dong et al., 2020; Barbarossa and Sardellitti, 2020; Chen et al., 2022; Roddenberry et al., 2021; Telyatnikov et al., 2023; Giusti et al., 2023a).
Figure 1: Workflow of TopoBenchmarkX, which consists of four main components, organized by their functionality: data modules, model modules, training modules, and communication modules.
TDL has evolved rapidly recently, providing novel algorithms for learning higher-order relational data supported on various topological domains, such as simplicial, cell, and combinatorial complexes. The works of Hajij et al. (2023) and Papillon et al. (2023b) have recently unified theory and notation of many TNNs. Moreover, the TAG ICML challenge 2023 (Papillon et al., 2023a) has contributed software implementations of topological message passing for most TNNs under TopoX (Hajij et al., 2024). Despite these theoretical and practical advancements, the rapid acceleration of TDL research poses challenges for the reproducibility and systematic comparative evaluation of TNNs (Papamarkou et al., 2024). In particular, open software, standardization, and benchmarking were recognized as key open challenges for TDL in a recent ICML 2024 position paper (Papamarkou et al., 2024). TopoBenchmarkX provides a robust framework to address these issues and ensure consistent and reliable progress in the field.
TopoBenchmarkX [^1] https://github.com/pyt-team/TopoBenchmarkX is an open-source and modular benchmarking framework for comparative evaluations of TNNs. Figure 1 visualizes the overall workflow and modularity of TopoBenchmarkX; see Section 4 for a more detailed outline. TopoBenchmarkX can generate topological datasets for benchmarking purposes, and carry out TNN comparisons based on various performance metrics. TopoBenchmarkX provides a high-level interface that reduces coding efforts, making it user-friendly. It can facilitate reproducibility of TNN research, and serves as a testbed to experiment with new ideas in TDL. Finally, TopoBenchmarkX is part of a larger suite of packages, TopoX (Hajij et al., 2024), and it is built with compatibility in mind, allowing integration with other tools and frameworks both within TopoX and in the broader PyTorch ecosystem.
Three aspects of TDL make benchmarking less straightforward: the scarcity of topological data, the non-trivial standardization of input and output across topological domains, and the diversity of TNN architectures. TopoBenchmarkX addresses these three limitations.
Generation of topological data. Although higher-order data arise in various physical systems, the complexity of capturing these interactions and the limitations of current experimental designs constrain data collection. As a result, higher-order data is often generated by lifting (i.e., transforming) graph datasets and equipping them with higher-order features. TopoBenchmarkX implements graph lifting algorithms to automate this generation process, thus mitigating the scarcity of topological datasets.
Standardization of input and output across topological domains. Each topological domain encodes higher-order relations differently. Further, each TNN architecture provides distinct mappings between inputs and outputs, and therefore processes higher-order data differently. TopoBenchmarkX standardizes the input-output pipeline through an interface that automatically manages the transition between topological domains and their associated TNNs.
Diversity of TNN architectures. There is a broad spectrum of TNNs in the TDL literature. Each TNN typically employs its own techniques to preprocess and encode data. This diversity complicates performance comparisons between models on different datasets. To address this issue, TopoBenchmarkX implements a pipeline for data preprocessing and predictive performance evaluation metrics. The pipeline is easily accessible to newcomers through a configuration file.
The rest of the paper is structured as follows. Section 2 outlines existing software for graph-based learning and benchmarking, as well as GDL. Section 3 introduces the notion of featured cell complex and formalizes the lifting of a graph to a cell complex. Section 4 introduces the main components of TopoBenchmarkX, and Section 5 demonstrates its usage via benchmarking experiments. The paper concludes with some final remarks, current limitations, and suggestions for future work in Section 6.
2 Existing software
Graph-based learning and GDL are supported by several software packages, including NetworkX (Hagberg et al., 2008), KarateClub (Rozemberczki et al., 2020), PyG (Fey and Lenssen, 2019), and DGL (Wang et al., 2019). NetworkX enables computations on graphs, while KarateClub implements unsupervised learning algorithms for graph-structured data. PyG and DGL provide functionality for GDL as well as graph-based learning.
Various software packages implement machine learning algorithms for higher-order domains such as HyperNetX (Liu et al., 2021), XGI (Landry et al., 2023), DHG (Feng et al., 2019), and TopoX (Hajij et al., 2024). HyperNetX facilitates computing with hypergraphs, while XGI enables computing with both hypergraphs and simplicial complexes. DHG implements deep learning algorithms for graphs and hypergraphs. TopoX is a suite of three packages: TopoNetX, TopoEmbedX, and TopoModelX, which together enable learning with TNNs. TopoNetX supports computations on graphs and higher-order domains, including simplicial complexes, cell complexes, path complexes, combinatorial complexes, and colored hypergraphs. TopoEmbedX implements algorithms to embed higher-order domains into Euclidean spaces. TopoModelX implements most of the TNNs surveyed in Papillon et al. (2023b).
The Open Graph Benchmark (OGB) is a set of datasets and a benchmark framework to facilitate reproducible graph machine learning research (Hu et al., 2020, 2021). The OGB is applicable to graph-based learning and GDL, but not to TDL. TopoBenchmarkX aims to fill this gap by providing algorithms that generate higher-order datasets and benchmark TNNs.
3 Featured topological domains and lifting
This section extends the concept of a featured graph to featured topological domains. To make the new definitions precise, cell complexes are used as a template for other topological domains. However, all definitions provided here are applicable to other complexes implemented in TopoBenchmarkX. This section also formalizes the lifting of a featured graph to a featured topological domain. See Appendix D for an extended introduction to other higher-order domains (e.g., simplicial complexes, hypergraphs).
3.1 Featured topological domains
A featured graph is a graph equipped with functions that map nodes or edges to features Sanchez-Lengeling et al. (2021).
Definition 1
Let $\mathcal{G}=(V,E)$ be a graph, with node set $V$ and edge set $E$. A featured graph is a tuple $\mathcal{G}_{F}=(V,E,F_{V},F_{E})$, where $F_{V}:V\to\mathbb{R}^{d}$ is a function that maps each node to a feature vector in $\mathbb{R}^{d}$ and $F_{E}:E\to\mathbb{R}^{k}$ is a function that maps each edge to a feature vector in $\mathbb{R}^{k}$.
Definition 2
A cell complex is a triplet $\mathcal{X}=(\mathcal{X}_{0},\mathcal{X}_{1},\mathcal{X}_{2})$ of finite ordered sets, where $\mathcal{X}_{0}$ is a set of nodes (or 00-cells), $\mathcal{X}_{1}$ is a set of edges (or $1$-cells) with each edge $e\in\mathcal{X}_{1}$ being an ordered pair of nodes $e=[v_{1},v_{2}]$ for $v_{1},v_{2}\in\mathcal{X}_{0}$, and $\mathcal{X}_{2}$ is a set of faces (or $2$-cells) with each face $\sigma\in\mathcal{X}_{2}$ being an ordered sequence of edges $\sigma=[e_{1},e_{2},\dots,e_{m(\sigma)}]$ forming a closed path without self-intersections and $m(\sigma)\geq 3$. A featured cell complex is a tuple $\mathcal{X}_{F}=(\mathcal{X}_{0},\mathcal{X}_{1},\mathcal{X}_{2},F_{\mathcal{X}_{0}},F_{\mathcal{X}_{1}},F_{\mathcal{X}_{2}}),$ where, for $0\leq i\leq 2$, $F_{\mathcal{X}_{i}}:\mathcal{X}_{i}\to\mathbb{R}^{k_{i}}$ maps each cell in $\mathcal{X}_{i}$ to a feature vector $\mathbb{R}^{k_{i}}$.
3.2 Liftings
In this paper, lifting refers to the process of mapping a featured graph to a featured topological domain through a well-defined procedure (Hajij et al., 2023). We provide a precise definition for the case of cell complexes. However, lifting to other topological domains can be defined similarly.
Definition 3
Let $\mathcal{G}_{F}=(V,E,F_{V},F_{E})$ be a featured graph. A lifting of $\mathcal{G}_{F}$ is a triplet $(\mathcal{C}_{F},\iota,L)$, where $\mathcal{C}_{F}=(\mathcal{X}_{0},\mathcal{X}_{1},\mathcal{X}_{2},F_{\mathcal{X}_{0}},F_{\mathcal{X}_{1}},F_{\mathcal{X}_{2}}),$ is a featured cell complex, $\iota$ is an embedding map that maps the nodes and edges of $\mathcal{G}$ to cells in $\mathcal{X}$, and $L$ is a collection of functions that map features of $\mathcal{G}_{F}$ to features of $\mathcal{C}_{F}$. The embedding map $\iota$ satisfies $\iota(v)=s_{v}$ for $v\in V$ with $s_{v}\in\mathcal{X}_{0}$ being a 00-cell corresponding to node $v$, and $\iota(e)=x_{e}$ for $e\in E$ with $x_{e}\in\mathcal{X}_{1}$ being a cell corresponding to edge $e$. The collection of functions in $L$ consists of a lifting procedure for the higher-order cells in $\mathcal{X}$ that are not part of $\mathcal{G}$ and a lifting procedure for the feature maps $\{F_{\mathcal{X}^{i}}\}_{i=0}^{2}$ based on $\mathcal{G}_{F}$.
Examples of liftings are given in Figure 2. Observe that a featured cell complex in Definition 3 acts as a template for a general topological domain to which the structural and feature information of the graph is lifted. In the general case, even if the graph features are empty (that is, if $F_{V}$ and $F_{E}$ are not defined), lifting still provides a way to map the graph structure to the cell complex structure.
Figure 2: An illustration of lifting a featured graph $\mathcal{G}_{F}$ (center) to two different topological domains: a simplicial complex (left) and a cell complex (right). The lifting from $\mathcal{G}_{F}$ (center) to the featured cell complex $\mathcal{C}_{F}$ (right) is represented by the triplet $(\mathcal{C}_{F},\iota,L)$, where $\iota$ is the embedding map that places the nodes and edges of $\mathcal{G}_{F}$ into the corresponding cells of $\mathcal{C}_{F}$, and $L$ is the collection of functions that maps the features of $\mathcal{G}_{F}$ to the features of $\mathcal{C}_{F}$.
Lifting maps can be fixed (Bodnar et al., 2021a; Hajij et al., 2023) or learnable (Battiloro et al., 2023a; Bernárdez et al., 2023). Fixed lifting maps are predefined and follow a rule-based procedure to lift a featured graph to a featured topological domain. For example, one could define a fixed lifting that always maps a node in the graph to a node in the topological domain, and an edge to a corresponding cell. In contrast, learnable lifting maps are determined using machine learning models that are trained to optimize the lifting process based on specific criteria or objective functions. During training, the model learns to embed the nodes and edges of the input graph into the cells of a topological domain in a way that optimizes the representation for downstream tasks, such as classification or clustering. When lifting a graph to a topological domain, both the features on the higher-order cells and the structure of the topological domain itself can be computed deterministically (Ramamurthy et al., 2023) or learned (Hajij et al., 2023). [^2] An unpooling operation from lower-rank cells, where features are given, to higher-rank cells, where features are to be lifted, is called a learnable feature lifter in Hajij et al. (2023).
4 Outline of TopoBenchmarkX library
TopoBenchmarkX implements a unified and flexible workflow that facilitates the addition of new datasets, transforms, and deep learning models, thus ensuring applicability across a wide range of tasks and enabling a broad cross-domain comparison, which is lacking in the TDL literature. Each module within TopoBenchmarkX is assigned a distinct role, while maintaining a consistent input-output structure, which provides a modular interface across all topological domains. Figure 1 outlines the TopoBenchmarkX modules, which are organized by functionality to data, model, training, and communication modules. Algorithm 1 exemplifies the TopoBenchmarkX execution pipeline via pseudo-code.
4.1 TopoBenchmarkX modules
Data modules. These modules include Loader, Transforms, PreProcessor, and Dataloader.
Loader. The loader module provides an interface for downloading and storing data; it includes the GraphLoader, HypergraphLoader, SimplicialLoader, and CellLoader classes. Their interfaces are built upon the widely-adopted InMemoryDataset, provided by PyG to enhance interoperability. The project webpage provides multiple in-depth tutorials on the library, including a step-by-step overview of incorporating customized data with these interfaces.
Transforms. Transforms modules are developed as sub-classes of the BaseTransform class provided by PyG. These modules are data manipulation, topology lifting, and feature lifting. The data manipulation module allows for performing any data transformation; e.g., the transformations available in PyG or TopoEmbedX can be easily adapted in TopoBenchmarkX. The topology lifting and feature lifting modules are responsible for lifting data from one topological domain to another. In particular, the former module relies on TopoNetX to lift the topology, while the latter utilizes the lifted topology to infer (e.g., project, concatenate) the features for higher-order topological structures. Each transform takes a Data object from PyG as input, performs the necessary computations, and outputs the modified Data object. The Transforms modules are composable operators that can be easily customized to perform various tasks.
Pre-processor. The PreProcessor class applies a sequence of transforms to a dataset. It takes a dataset object and a sequence of transforms as input, and iterates over the dataset, applying a set of transforms sequentially to each instance in the dataset. To avoid re-computing the same transforms multiple times, the pre-processed dataset is automatically saved in a unique folder for each transform configuration. This ensures that each dataset needs to be processed only once per configuration, addressing the potential time-consuming nature of preprocessing large datasets. PreProcessor also generates and/or loads data splits according to the chosen strategy; for example, random splits with predefined proportions, k-fold cross-validation, or fixed splits.
Dataloader. The dataloader module provides a consistent interface for batch training across all topological domains. TopoBenchmarkX supports mini-batching for graphs, hypergraphs, simplicial complexes, and cell complexes.
Model modules. Neural networks modules form the core of the modeling pipeline. The encoder module maps initial data features into a latent space and applies a learnable transformation before passing the data to the TNN model. The backbone TNN can be imported from TopoModelX, PyG, or it can be built on a custom basis, allowing flexibility in model selection (see Table 6 in Appendix E.2). The wrapper ensures that the correct input is provided to the forward pass of the model and collects the output in a dictionary. The wrapper plays a crucial role in the integration of any model in TopoBenchmarkX, making it possible to use existing and newly developed models. The readout module processes latent representations of the neural network into final output predictions. The Loss module is responsible for defining the loss function from the PyTorch library. The optimizer module defines the optimizer and scheduler. It allows for easy integration of any optimizer and scheduler from torch.optim, which provides flexibility in optimizing model training. The evaluator module is responsible for computing metrics during training and inference. It is built upon torchmetrics and supports standard metrics for classification and regression tasks.
Training and communication modules. The Model class defines a training pipeline for all domains (lines 14-19 of Algorithm 1). It inherits from LightningModule, and requires the Encoder, Wrapper, Backbone, Readout, Evaluator, Loss, and Optimizer objects as inputs. The lightning.Trainer automates the training, evaluation, and testing processes. Extra functionalities can be added using callbacks modules, and users can monitor the training process using different loggers (e.g. wandb, tensorboard). Both are common practice in Lightning, and they are referred to as communication modules in TopoBenchmarkX.
Algorithm 1 Execution pipeline for model training in TopoBenchmarkX
1:  Input: General cfg configuration file
2:  dataset $\leftarrow$ Loader(cfg.dataset) # Dataset loading
3:  splits $\leftarrow$ PreProcessor(dataset, cfg.transforms) # Transforms and splits
4:  dataloader $\leftarrow$ Dataloader(dataset) # Batch generator
5:  model $\leftarrow$ Model( # Model initialization
6:      nn.Encoder(cfg.model),
7:      nn.Backbone(cfg.model),
8:      nn.BackboneWrapper(cfg.model),
9:      nn.Readout(cfg.model),
10:      *[Evaluator(cfg.evaluator), Optimizer(cfg.optimizer), Loss(cfg.loss)])
11:  trainer $\leftarrow$ lightning.Trainer(cfg.trainer, cfg.callbacks, cfg.logger)
12:  Model training:
13:  trainer.fit(model, dataloader) # Model training
14:  Model step for each batch:
15:  batch $\leftarrow$ self.encoder(batch) # Feature encoder
16:  model_out $\leftarrow$ self.forward(batch) # TNN
17:  model_out $\leftarrow$ self.readout(model_out, batch) # Readout
18:  model_out $\leftarrow$ self.loss(model_out, batch) # Loss computation
19:  self.evaluator.update(model_out) # Evaluator update
4.2 Topological liftings
This section explores different (structural) liftings implemented in TopoBenchmarkX. In particular, liftings of graphs to cell complexes, simplicial complexes, and hypergraphs are discussed. [^3] Note that the modular design of TopoBenchmarkX easily allows for the addition of new lifting transforms between any pair of topological domains.
From graphs to cell complexes: cycle-based liftings. A graph is lifted to a cell complex as follows. First, a finite set of cycles (closed loops) within the graph is identified. Second, each identified cycle in the graph is associated with a 2-cell, such that the boundary of the 2-cell is the cycle. The nodes and edges of the cell complex are inherited from the graph.
From graphs to simplicial complexes: clique complexes. By lifting a graph to a simplicial complex, both pairwise and higher-order interactions are captured. For a graph, the corresponding clique complex is formed by considering every complete subgraph (clique) as a simplex. Specifically, each node is a 0-simplex, each edge (clique of size 2) is a 1-simplex, each triangle (clique of size 3) is a 2-simplex, and so forth. In general, a clique of size $k+1$ becomes a $k$-simplex.
From graphs to simplicial complexes: neighbor complexes. Neighbor complexes lift the neighborhoods of nodes to become simplices as follows. For each node in the graph, take the set of its neighbors and the node itself. These sets are then treated as simplices. The dimension of each simplex depends on the degree of the nodes. For example, a node with $d$ neighbors forms a $d$-simplex.
From graphs to hypergraphs: $k$-hop liftings. Let $\mathcal{G}=(V,E)$ be a graph and $\mathcal{H}=(V,\mathcal{E})$ be a hypergraph. The $k$-neighborhood $N_{k}(v)$ of a node $v\in V$ in $\mathcal{G}$ consists of all nodes reachable within $k$ steps from $v$. To lift $\mathcal{G}$ to $\mathcal{H}$, for each node $v\in V$, determine $N_{k}(v)$ and create a hyperedge $e_{v}$ in $\mathcal{H}$ such that $e_{v}=N_{k}(v)$. Thus, the set of hyperedges in $\mathcal{H}$ is given by =$\mathcal{E}=\{N_{k}(v)\mid v\in V\}$.
5 Numerical experiments
This section presents numerical experiments that exemplify the wide range of functionality of TopoBenchmarkX, and provide a first attempt to perform a broad cross-domain comparison. First, the general experimental setup is described. The experiments compare twelve graph, hypergraph and topological neural network models across four learning tasks related to twenty-two datasets. Subsequently, the main benchmark results are presented. Finally, it is shown how TopoBenchmarkX facilitates comparisons in TDL by conducting an ablation study for different types of signal propagation [^4] See Appendix E for an introduction to TNNs and higher-order message passing on topological domains..
5.1 Setup
Learning tasks and datasets. Four learning tasks are carried out, namely node classification (seven datasets), node regression (seven datasets), graph classification (seven datasets), and graph regression (one dataset). For node classification, cocitation datasets (Cora, Citeseer, and Pubmed), as well as heterophilic datasets (Amazon rating, Minesweeper, Roman empire, and Tolokers) are used (Platonov et al., 2023). For node regression, the election, bachelor, birth, death, income, migration, and unemployment datasets from US election map networks are adapted (Jia and Benson, 2020). In these datasets, nodes represent US states, edges connect neighboring states, and each state is characterized by demographic and election statistics. For each dataset, one statistic is set to be the output variable, while the remaining statistics serve as node features, referring to each such dataset with the output statistic’s name. For graph classication, the TUDataset is used, consisting of the MUTAG, PROTEINS, NCI1, NCI109, IMDB-BIN, IMDB-MUL, and REDDIT datasets; see Morris et al. (2020). For graph regression, the ZINC dataset is employed (Irwin et al., 2012). See Table 7 in Appendix F for descriptive summaries of the datasets.
Models. The twelve neural network models are supported on four domains: graphs, hypergraphs, simplicial complexes, and cell complexes. In particular, three GNNs (GCN, GIN, and GAT), three hypergraph neural networks (EDGNN, AllSetTransformer, and UniGNN2), three simplicial neural networks (SCN, SCCN, and SCCNN), and three cell complex neural networks (CCXN, CWN, and CCCN) are used in the experiments. Please see Appendix E.2 for further details about these architectures, including a comparison of their hyperparameter budget.
Training and evaluation. Five splits are generated for each dataset. Each split comprises 50%/25%/25% of data points in the training, validation, and test set, respectively. As an exception, the pre-defined splits for the ZINC dataset are used (Irwin et al., 2012). The optimal hyperparameter set is selected based on the best average performance over five validation splits. See Appendix C.2 for more details on the hyperparameter search. One performance metric per dataset is reported. More specifically, the predictive accuuracy is reported for Cora, Citeseer, Pubmed, Amazon, Empire, MUTAG, PROTEINS, NCI1, NCI109, IMDB-BIN, IMDB-MUL, and REDDIT; the AUC-ROC metric is reported for Minesweeper and Tolokers; the mean-squared error is reported for election, bachelor, birth, death, income, migration, and unemployment; and the mean-absolute error is reported for ZINC. For each performance metric, its mean and standard deviation are computed over the five test sets per dataset and are reported in Table 1 (note that OOM stands for ‘out of memory’ in Table 1).
5.2 Main results
As seen from Table 1, higher-order neural networks (based on hypergraphs, simplicial and cell complexes) achieve the best performance on sixteen of twenty-two datasets, whereas GNNs achieve the best performance on seven datasets. GNNs perform best on node regression in the majority of cases (five out of seven). These best results obtained by GNNs are closely matched by TNNs, since the latter attain metrics within one standard deviation from the former. In contrast, in nine out of sixteen datasets TNNs outperform GNNs, and attain performance metrics that are higher by more than one standard deviation with respect to GNNs. In other words, in the cases where higher-order networks outperform GNNs, the performance gap is more pronounced. It is also noted that, for demonstration purposes, only one fixed lifting is considered to transform graph data to each of the considered topological domains (see Appendix C.3). These results suggest that TNNs hold an advantage over GNNs, even without lifting optimization. Most importantly for the context of this paper, the benchmarks demonstrate the extent of comparisons that can be performed with TopoBenchmarkX across models and datasets.
Table 1: Cross-domain comparison: results are shown as mean and standard deviation. The best result is bold and shaded in grey, while those within one standard deviation are in blue-shaded boxes.
Dataset GCN GIN GAT AST EDGNN UniGNN2 CWN CCCN SCCNN SCN Node-level tasks Cora 87.09 ± 0.20 87.21 ± 1.89 86.71 ± 0.95 88.92 ± 0.44 87.06 ± 1.09 86.97 ± 0.88 86.32 ± 1.38 87.44 ± 1.28 82.19 ± 1.07 82.27 ± 1.34 Citeseer 75.53 ± 1.27 73.73 ± 1.23 74.41 ± 1.75 73.85 ± 2.21 74.93 ± 1.39 74.72 ± 1.08 75.20 ± 1.82 75.63 ± 1.58 70.23 ± 2.69 71.24 ± 1.68 Pubmed 89.40 ± 0.30 89.29 ± 0.41 89.44 ± 0.24 89.62 ± 0.25 89.04 ± 0.51 89.34 ± 0.45 88.64 ± 0.36 88.52 ± 0.44 88.18 ± 0.32 88.72 ± 0.50 Amazon 49.56 ± 0.55 49.16 ± 1.02 50.17 ± 0.59 50.50 ± 0.27 48.18 ± 0.09 49.06 ± 0.08 51.90 ± 0.15 50.26 ± 0.17 OOM OOM Empire 78.16 ± 0.32 79.56 ± 0.20 84.02 ± 0.51 79.50 ± 0.13 81.01 ± 0.24 77.06 ± 0.20 81.81 ± 0.62 82.14 ± 0.00 89.15 ± 0.32 88.79 ± 0.46 Minesweeper 87.52 ± 0.42 87.82 ± 0.34 89.64 ± 0.43 81.14 ± 0.05 84.52 ± 0.05 78.02 ± 0.00 88.62 ± 0.04 89.42 ± 0.00 89.0 ± 0.00 90.32 ± 0.11 Tolokers 83.02 ± 0.71 80.72 ± 1.19 84.43 ± 1.00 83.26 ± 0.10 77.53 ± 0.01 77.35 ± 0.03 OOM OOM OOM OOM Election 0.31 ± 0.02 0.28 ± 0.02 0.29 ± 0.02 0.29 ± 0.01 0.34 ± 0.02 0.37 ± 0.02 0.34 ± 0.02 0.31 ± 0.02 0.51 ± 0.03 0.46 ± 0.04 Bachelor 0.29 ± 0.02 0.31 ± 0.03 0.28 ± 0.02 0.30 ± 0.03 0.29 ± 0.02 0.31 ± 0.02 0.33 ± 0.03 0.31 ± 0.02 0.34 ± 0.03 0.32 ± 0.02 Birth 0.72 ± 0.09 0.72 ± 0.09 0.71 ± 0.09 0.71 ± 0.08 0.70 ± 0.07 0.73 ± 0.10 0.72 ± 0.09 0.71 ± 0.09 0.79 ± 0.12 0.71 ± 0.08 Death 0.51 ± 0.04 0.52 ± 0.04 0.51 ± 0.04 0.49 ± 0.05 0.52 ± 0.05 0.51 ± 0.05 0.54 ± 0.06 0.54 ± 0.06 0.55 ± 0.05 0.52 ± 0.05 Income 0.22 ± 0.03 0.21 ± 0.02 0.20 ± 0.02 0.21 ± 0.02 0.23 ± 0.02 0.23 ± 0.02 0.25 ± 0.03 0.23 ± 0.02 0.28 ± 0.03 0.25 ± 0.02 Migration 0.80 ± 0.12 0.80 ± 0.10 0.77 ± 0.13 0.78 ± 0.12 0.80 ± 0.12 0.79 ± 0.12 0.84 ± 0.13 0.84 ± 0.12 0.90 ± 0.14 0.92 ± 0.20 Unempl 0.25 ± 0.03 0.22 ± 0.02 0.23 ± 0.03 0.22 ± 0.02 0.26 ± 0.03 0.28 ± 0.02 0.25 ± 0.03 0.24 ± 0.03 0.43 ± 0.04 0.38 ± 0.04 Graph-level tasks MUTAG 69.79 ± 6.80 79.57 ± 6.13 72.77 ± 2.77 71.06 ± 6.49 80.00 ± 4.90 80.43 ± 4.09 80.43 ± 1.78 77.02 ± 9.32 76.17 ± 6.63 73.62 ± 6.13 PROTEINS 75.70 ± 2.14 75.20 ± 3.30 76.34 ± 1.66 76.63 ± 1.74 73.91 ± 4.39 75.20 ± 2.96 76.13 ± 2.70 73.33 ± 2.30 74.19 ± 2.86 75.27 ± 2.14 NCI1 72.86 ± 0.69 74.26 ± 0.96 75.00 ± 0.99 75.18 ± 1.24 73.97 ± 0.82 73.02 ± 0.92 73.93 ± 1.87 76.67 ± 1.48 76.60 ± 1.75 74.49 ± 1.03 NCI109 72.20 ± 1.22 74.42 ± 0.70 73.80 ± 0.73 73.75 ± 1.09 74.93 ± 2.50 70.76 ± 1.11 73.80 ± 2.06 75.35 ± 1.50 77.12 ± 1.07 75.70 ± 1.04 IMDB-BIN 72.00 ± 2.48 70.96 ± 1.93 69.76 ± 2.65 70.32 ± 3.27 69.12 ± 2.92 71.04 ± 1.31 70.40 ± 2.02 69.12 ± 2.82 70.88 ± 2.25 70.80 ± 2.38 IMDB-MUL 49.97 ± 2.16 47.68 ± 4.21 50.13 ± 3.87 50.51 ± 2.92 49.17 ± 4.35 49.76 ± 3.55 49.71 ± 2.83 47.79 ± 3.45 48.75 ± 3.98 49.49 ± 5.08 REDDIT 76.24 ± 0.54 81.96 ± 1.36 75.68 ± 1.00 74.84 ± 2.68 83.24 ± 1.45 75.56 ± 3.19 85.52 ± 1.38 85.12 ± 1.29 77.24 ± 1.87 71.28 ± 2.06 ZINC 0.62 ± 0.01 0.57 ± 0.04 0.61 ± 0.01 0.59 ± 0.02 0.51 ± 0.01 0.60 ± 0.01 0.34 ± 0.01 0.34 ± 0.02 0.36 ± 0.02 0.53 ± 0.04
5.3 Ablation study
Graph and hypergraph (neural network) models differ from simplicial and cell complex (neural network) models in terms of the domains and representations they support. Graph and hypergraph models can output two types of representations: node representations and edge or hyperedge representations. In contrast, the output of simplicial and cell models depends on the different types of cells present (00-cell up to $n$-cells) and the model itself. For example, a simplicial or cell complex model may process an $n$-cell input but may not produce an $n$-cell output. The backbone_wrapper in TopoBenchmarkX addresses these differences in the underlying domains of the models.
There is a second difference, which is inherent in the TNNs themselves. Consider a downstream classification task. For graphs, the standard practice is to perform classification over pooled node features. However, this aspect has not been extensively studied in the TDL literature. For instance, a simplicial or cell model may update 1-cell representations (edges) or 2-cell representations (cycles or triangles) while leaving 0-cell (nodes) representations unchanged, making direct pooling over nodes potentially meaningless. One can consider more complicated update processes in which different $n$-cell representations are combined, which makes the notion of pooling more intricate for higher-order domains, in general. These architectural implications are complex and are related to open research questions in TDL.
However, to compare different neural network architectures, the second aforementioned difference must be addressed. To this end, the ablation study of this section considers two types of readouts to perform a rigorous evaluation: direct readout (DR), in which the downstream task is performed directly over the 00-cell representation, and signal down-propagation (SDP), in which information from higher-order cell representations is iteratively fused down to 00-cell representations using appropriate incidence matrices, followed by a linear projection over the concatenated $n-1$-cell signal and the fused $n-1$-cell representation. For instance, if a simplicial or cell complex model outputs 00-cell, $1$-cell, and $2$-cell representations, the signal propagates from $2$-cells to $1$-cells, and from $1$-cells to 00-cells during readout. The downstream task is then performed over the 00-cell representations. [^5] See Appendix E for an introduction to TNNs and higher-order message passing on topological domains.
Table 2 shows that the best-performing readout type depends on the model’s interior signal propagation. The CWN model does not update the 00-cell representation during propagation and the SDP readout strategy performs notably better, as expected. Conversely, CCCN, SCCNN, and SCN propagate information to 00-cells, so incorporating SDP readout leads to either small or minor variations in performance metrics. Please refer to Appendix C.3 for more detailed results.
These results highlight the importance of considering the structural properties of TNNs. The performance variations observed in this ablation study emphasize how critical architectural and lifting decisions are in higher-order learning models. By facilitating comparisons across a wide range of models and datasets, TopoBenchmarkX enables researchers to derive new insights and advance TDL.
Table 2: An ablation study compares the performance of CWN, CCCN, SCCNN, and SCN models on various datasets using two readout strategies, direct readout (DR) and signal down-propagation (SDP). SDP generally enhances CWN performance, whereas the effect of SDP on CCCN, SCCNN, and SCN varies based on their internal signal propagation mechanisms. Means and standard deviations of performance metrics are shown. The best results are shown in bold for each model and readout type.
Dataset CWN CCCN SCCNN SCN DR SDP DR SDP DR SDP DR SDP Node-level tasks Cora 74.95 ± 0.98 86.32 ± 1.38 87.44 ± 1.28 87.68 ± 1.17 82.19 ± 1.07 80.65 ± 2.39 82.27 ± 1.34 79.91 ± 1.18 Citeseer 70.49 ± 2.85 75.20 ± 1.82 75.63 ± 1.58 74.91 ± 1.25 70.23 ± 2.69 69.03 ± 2.01 71.24 ± 1.68 70.40 ± 1.53 Pubmed 86.94 ± 0.68 88.64 ± 0.36 88.52 ± 0.44 88.67 ± 0.39 88.18 ± 0.32 87.78 ± 0.58 88.72 ± 0.50 88.62 ± 0.44 Amazon 45.58 ± 0.25 51.90 ± 0.15 50.55 ± 0.15 50.26 ± 0.17 OOM OOM OOM OOM Empire 66.13 ± 0.03 81.81 ± 0.62 82.14 ± 0.00 82.51 ± 0.00 89.15 ± 0.32 88.73 ± 0.12 85.89 ± 0.34 88.79 ± 0.46 Minesweeper 48.82 ± 0.00 88.62 ± 0.04 89.42 ± 0.00 89.85 ± 0.00 87.40 ± 0.00 89.00 ± 0.00 90.32 ± 0.11 90.27 ± 0.36 Election 0.60 ± 0.04 0.34 ± 0.02 0.31 ± 0.02 0.31 ± 0.01 0.51 ± 0.03 0.56 ± 0.04 0.46 ± 0.04 0.51 ± 0.03 Bachelor 0.33 ± 0.03 0.33 ± 0.03 0.32 ± 0.02 0.31 ± 0.02 0.34 ± 0.03 0.34 ± 0.03 0.32 ± 0.02 0.32 ± 0.03 Birth 0.81 ± 0.11 0.72 ± 0.09 0.71 ± 0.09 0.72 ± 0.05 0.79 ± 0.12 0.83 ± 0.12 0.71 ± 0.08 0.80 ± 0.11 Death 0.55 ± 0.05 0.54 ± 0.06 0.54 ± 0.06 0.54 ± 0.06 0.55 ± 0.05 0.58 ± 0.05 0.52 ± 0.05 0.56 ± 0.05 Income 0.36 ± 0.04 0.25 ± 0.03 0.23 ± 0.02 0.23 ± 0.02 0.28 ± 0.03 0.31 ± 0.03 0.25 ± 0.02 0.27 ± 0.02 Migration 0.90 ± 0.16 0.84 ± 0.13 0.84 ± 0.10 0.84 ± 0.12 0.90 ± 0.14 0.93 ± 0.17 0.92 ± 0.20 0.96 ± 0.23 Unempl 0.46 ± 0.04 0.25 ± 0.03 0.24 ± 0.03 0.25 ± 0.03 0.43 ± 0.04 0.45 ± 0.04 0.38 ± 0.04 0.41 ± 0.03 Graph-level tasks MUTAG 69.68 ± 8.58 80.43 ± 1.78 80.85 ± 5.42 77.02 ± 9.32 76.17 ± 6.63 70.64 ± 3.16 71.49 ± 2.43 73.62 ± 6.13 PROTEINS 76.13 ± 1.80 76.13 ± 2.70 73.55 ± 3.43 73.33 ± 2.30 74.19 ± 2.86 74.98 ± 1.92 75.27 ± 2.14 74.77 ± 1.69 NCI1 68.52 ± 0.51 73.93 ± 1.87 76.67 ± 1.48 77.65 ± 1.28 76.60 ± 1.75 75.60 ± 2.45 75.27 ± 1.57 74.49 ± 1.03 NCI109 68.19 ± 0.65 73.80 ± 2.06 75.35 ± 1.50 74.83 ± 1.18 77.12 ± 1.07 75.43 ± 1.94 74.58 ± 1.29 75.70 ± 1.04 IMDB-BIN 70.40 ± 2.02 69.28 ± 2.57 69.12 ± 2.82 69.44 ± 2.46 70.88 ± 2.25 69.28 ± 5.69 70.80 ± 2.38 68.64 ± 3.90 IMDB-MUL 49.71 ± 2.83 49.87 ± 2.33 49.01 ± 2.63 47.79 ± 3.45 48.75 ± 3.98 46.67 ± 3.13 48.16 ± 2.89 49.49 ± 5.08 REDDIT 76.20 ± 0.86 85.52 ± 1.38 85.12 ± 1.29 83.32 ± 0.73 75.56 ± 3.46 77.24 ± 1.87 71.28 ± 2.06 69.68 ± 4.00 ZINC 0.70 ± 0.00 0.34 ± 0.01 0.35 ± 0.02 0.34 ± 0.02 0.36 ± 0.01 0.36 ± 0.02 0.59 ± 0.01 0.53 ± 0.04
6 Concluding remarks, limitations, and future work
This paper has introduced TopoBenchmarkX, an open-source benchmarking framework for TDL. TopoBenchmarkX simplifies the benchmarking process and accelerates research by organizing the TDL pipeline into a sequence of modular steps. A key feature of TopoBenchmarkX is its ability to lift between topological domains, allowing richer data representations and more detailed analyses. This capability supports the mapping of graph topology and features to higher-order topological domains, such as simplicial and cell complexes. The effectiveness and applicability of TopoBenchmarkX has been illustrated by benchmarking several TDL architectures on many learning tasks and datasets, gaining some preliminary insights about the relative advantages of different models.
The modular and scalable properties of TopoBenchmarkX open up several opportunities for future enhancements, and simultaneously allow us to address its current limitations. One area for future work is the integration of learnable liftings, not yet implemented in the framework. This is a promising direction, since learnable liftings can learn optimal representations of topological domains for specific downstream learning tasks. A second relevant limitation is the current lack of higher-order standard datasets. Thus, a second future work direction is to incorporate higher-order datasets into TopoBenchmarkX, giving users the option to choose between downloading built-in datasets (not yet available software feature) and generating synthetic higher-order data by lifting (already available software feature). Lastly, a third critical aspect pertains to the set of performance metrics. A wider range of alternative evaluation metrics specific to TDL needs to be developed and implemented to assess expressivity, explainability, and fairness (Papamarkou et al., 2024). More generally, users and researchers are encouraged to contribute to the development of TopoBenchmarkX in any of these three or other preferred directions.
Appendix A Societal impact and ethics
TopoBenchmarkX aims to standardize benchmarking and accelerate research in Topological Deep Learning, so we do not expect it to have any direct negative societal impact from its usage. Moreover, authors and contributors advocate for responsible use of TopoBenchmarkX, adherence to ethical guidelines, and promote transparency, fairness, and inclusivity in research.
The TopoBenchmarkX library will be constantly maintained to respect proprietary content. It will implement strict revision processes to ensure that all code implementations, libraries, and datasets have open-source licenses that guarantee their legitimate usage within the framework.
Appendix B Code availability and reproducibility
The code for TopoBenchmarkX is available on GitHub under MIT license: https://github.com/pyt-team/TopoBenchmarkX. The codebase adheres to best software practices, employs continuous integration, is fully documented, and provides API documentation on its website https://pyt-team.github.io/topobenchmarkx/index.html.
The README.md comprehensively covers all aspects of library installation and development. To replicate the experiments outlined in this work, refer to the ‘Experiments Reproducibility’ section in the README.md. Additionally, the ‘tutorials’ folder contains various tutorials showcasing the integration of new models, datasets, and transforms.
Appendix C Training details and additional results
In this section, we delve into the details of our hyperparameter search methodology, optimization strategy, computational resources utilized for conducting the experiments, and present additional results and analyses.
C.1 Experiments’ configuration and model execution
To automate the configuration of the TopoBenchmarkX modules we utilize hydra package [Yadan, 2019]. In particular, we rely on hierarchical configuration groups and registers to facilitate the easy use of the library. Specifically, there is no need to meticulously select each module to execute the pipeline in any of the domains. Simply choosing a dataset and a model automatically configures a full default pipeline, eliminating the need for manual intervention.
TopoBenchmarkX also automates the model execution and learning through the lightning library [Lisa and Bot, 2017], orchestrating the training, validation and testing processes –as well as handling any logging and callbacks.
C.2 Hyperaparameter search
Five splits are generated for each dataset to ensure a fair evaluation of the models across domains. Each split comprises 50% training data, 25% validation data, and 25% test data. An exception is made for the ZINC dataset, where pre-defined splits are used [Irwin et al., 2012].
Each model in every domain has numerous specific hyperparameters that can be tuned to enhance performance. TNNs, in particular, have multiple additional parameters that can potentially improve results. To avoid the combinatorial explosion of possible hyperparameter sets, the search space is limited to the hyperparameters common across all models. The grid-search strategy is employed to find the optimal hyperparameters for each model and dataset. Specifically, the encoder hidden dimension is varied over [32, 64, 128], the encoder dropout is varied over [0.25, 0.5], the number of backbone layers is varied over [1, 2, 3, 4], the learning rate is varied over [0.01, 0.001], and the batch size is varied over [128, 256]. Additionally, for the cellular and simplicial domains, the readout type is varied over [direct readout (DR), signal down-propagation (SDP)]. If the model cannot fit into memory, the batch size, encoder hidden dimension, and number of backbone layers are decreased until the model training fits into the GPU memory.
For node-level task datasets, validation is conducted after each training epoch, continuing until either the maximum number of epochs is reached or the optimization metric fails to improve for 50 consecutive validation epochs. The minimum number of epochs is set to 50. Conversely, for graph-level tasks, validation is performed every 5 training epochs, with training halting if the performance metric does not improve on the validation set for the last 10 validation epochs. To optimize the models, torch.optim.Adam is combined with torch.optim.lr_scheduler.StepLR wherein the step size was set to 50 and the gamma value to 0.5. In total more than 100,000 runs have been performed to obtain the results. The optimal hyperparameter set is generally selected based on the best average performance over five validation splits. For the ZINC dataset, five different initialization seeds are used to obtain the average performance.
The hyperparameter search is executed on a Linux machine with 256 cores, 1TB of system memory, and 8 NVIDIA A30 GPUs, each with 24GB of GPU memory.
C.3 Results
Table 3 additionally presents the results for the CCXN and SCCN networks, which on average, perform slightly worse than the other models. As shown in Table 4, the CCXN network performs better when using the SDP readout strategy, although the performance improvements are not as significant as those achieved by the CWN network with the same strategy. The SCCN model benefits more from utilizing the SDP readout compared to other simplicial domain models (SCN and SCCNN), showing performance improvements in 9 out of 21 cases, while SCCNN and SCN show improvements in only 3 and 5 cases, respectively. Overall, cellular models demonstrate performance enhancements on 15, 19, and 8 datasets for CCXN, CWN, and CCCN, respectively, with the SDP readout strategy. Conversely, simplicial domain models achieve 9, 3, and 5 performance improvements for SCCN, SCCNN, and SCN, respectively, with the same SDP strategy. Based on this analysis and the findings in Section 5.3, it is recommended for the new people in the TDL to opt for the SDP readout strategy if they prefer not to delve into the intricate details of interior propagation in cellular and simplicial models.
Note that for demonstration purposes, only one fixed lifting is considered for lifting graphs to each of the considered topological domains, leaving out of the scope of this paper the analysis of the optimal lifting in each particular scenario. [^6] Note that learnable liftings can further optimize the predictive capacity of higher-order networks. In particular, the clique complex lifting is applied to lift the graph to a simplicial domain. Cycle-based lifting is employed to lift the graph into the cellular domain, and $k$-hop lifting is utilized to lift the graph into a hypergraph, where $k=1$. Projection lifting is applied to lift the features, specifically, the features of the $n-1$-cell projected to the $n$-cell by multiplying $n-1$-cell representation to corresponding incidence matrices.
Table 3: Cross-domain comparison: results are shown as mean and standard deviation. The best result is bold and shaded in grey, while those within one standard deviation are in blue-shaded boxes.
Node-level tasks Dataset GCN GIN GAT AST EDGNN UniGNN2 CCXN CWN CCCN SCCN SCCNN SCN Cora 87.09 ± 0.2 87.21 ± 1.89 86.71 ± 0.95 88.92 ± 0.44 87.06 ± 1.09 86.97 ± 0.88 86.79 ± 1.81 86.32 ± 1.38 87.44 ± 1.28 80.86 ± 2.16 82.19 ± 1.07 82.27 ± 1.34 Citeseer 75.53 ± 1.27 73.73 ± 1.23 74.41 ± 1.75 73.85 ± 2.21 74.93 ± 1.39 74.72 ± 1.08 74.67 ± 2.24 75.2 ± 1.82 75.63 ± 1.58 69.6 ± 1.83 70.23 ± 2.69 71.24 ± 1.68 Pubmed 89.4 ± 0.3 89.29 ± 0.41 89.44 ± 0.24 89.62 ± 0.25 89.04 ± 0.51 89.34 ± 0.45 88.91 ± 0.47 88.64 ± 0.36 88.52 ± 0.44 88.37 ± 0.48 88.18 ± 0.32 88.72 ± 0.5 Amazon 49.56 ± 0.55 49.16 ± 1.02 50.17 ± 0.59 50.5 ± 0.27 48.18 ± 0.09 49.06 ± 0.08 48.93 ± 0.14 51.9 ± 0.15 50.26 ± 0.17 OOM OOM OOM Empire 78.16 ± 0.32 79.56 ± 0.2 84.02 ± 0.51 79.5 ± 0.13 81.01 ± 0.24 77.06 ± 0.2 81.44 ± 0.31 81.81 ± 0.62 82.14 ± 0.0 88.27 ± 0.14 89.15 ± 0.32 88.79 ± 0.46 Minesweeper 87.52 ± 0.42 87.82 ± 0.34 89.64 ± 0.43 81.14 ± 0.05 84.52 ± 0.05 78.02 ± 0.0 $8$ 0.36 88.62 ± 0.04 89.42 ± 0.0 89.07 ± 0.25 89.0 ± 0.0 90.32 ± 0.11 Tolokers 83.02 ± 0.71 80.72 ± 1.19 84.43 ± 1.0 83.26 ± 0.1 77.53 ± 0.01 77.35 ± 0.03 OOM OOM OOM OOM OOM OOM Election 0.31 ± 0.02 0.28 ± 0.02 0.29 ± 0.02 0.29 ± 0.01 0.34 ± 0.02 0.37 ± 0.02 0.35 ± 0.02 0.34 ± 0.02 0.31 ± 0.02 0.53 ± 0.03 0.51 ± 0.03 0.46 ± 0.04 Bachelor 0.29 ± 0.02 0.31 ± 0.03 0.28 ± 0.02 0.3 ± 0.03 0.29 ± 0.02 0.31 ± 0.02 0.32 ± 0.03 0.33 ± 0.03 0.31 ± 0.02 0.36 ± 0.02 0.34 ± 0.03 0.32 ± 0.02 Birth 0.72 ± 0.09 0.72 ± 0.09 0.71 ± 0.09 0.71 ± 0.08 0.7 ± 0.07 0.73 ± 0.1 0.74 ± 0.11 0.72 ± 0.09 0.71 ± 0.09 0.82 ± 0.09 0.79 ± 0.12 0.71 ± 0.08 Death 0.51 ± 0.04 0.52 ± 0.04 0.51 ± 0.04 0.49 ± 0.05 0.52 ± 0.05 0.51 ± 0.05 0.54 ± 0.06 0.54 ± 0.06 0.54 ± 0.06 0.58 ± 0.06 0.55 ± 0.05 0.52 ± 0.05 Income 0.22 ± 0.03 0.21 ± 0.02 0.2 ± 0.02 0.21 ± 0.02 0.23 ± 0.02 0.23 ± 0.02 0.25 ± 0.03 0.25 ± 0.03 0.23 ± 0.02 0.29 ± 0.03 0.28 ± 0.03 0.25 ± 0.02 Migration 0.8 ± 0.12 0.8 ± 0.1 0.77 ± 0.13 0.78 ± 0.12 0.8 ± 0.12 0.79 ± 0.12 0.85 ± 0.18 0.84 ± 0.13 0.84 ± 0.12 0.91 ± 0.18 0.9 ± 0.14 0.92 ± 0.2 Unempl 0.25 ± 0.03 0.22 ± 0.02 0.23 ± 0.03 0.22 ± 0.02 0.26 ± 0.03 0.28 ± 0.02 0.27 ± 0.03 0.25 ± 0.03 0.24 ± 0.03 0.43 ± 0.04 0.43 ± 0.04 0.38 ± 0.04 Graph-level tasks MUTAG 69.79 ± 6.8 79.57 ± 6.13 72.77 ± 2.77 71.06 ± 6.49 80.0 ± 4.9 80.43 ± 4.09 74.89 ± 5.51 80.43 ± 1.78 77.02 ± 9.32 70.64 ± 5.9 76.17 ± 6.63 73.62 ± 6.13 PROTEINS 75.7 ± 2.14 75.2 ± 3.3 76.34 ± 1.66 76.63 ± 1.74 73.91 ± 4.39 75.2 ± 2.96 75.63 ± 2.57 76.13 ± 2.7 73.33 ± 2.3 75.05 ± 2.76 74.19 ± 2.86 75.27 ± 2.14 NCI1 72.86 ± 0.69 74.26 ± 0.96 75.0 ± 0.99 75.18 ± 1.24 73.97 ± 0.82 73.02 ± 0.92 74.86 ± 0.82 73.93 ± 1.87 76.67 ± 1.48 76.17 ± 1.39 76.6 ± 1.75 74.49 ± 1.03 NCI109 72.2 ± 1.22 74.42 ± 0.7 73.8 ± 0.73 73.75 ± 1.09 74.93 ± 2.5 70.76 ± 1.11 75.66 ± 1.3 73.8 ± 2.06 75.35 ± 1.5 75.49 ± 1.39 77.12 ± 1.07 75.7 ± 1.04 IMDB-BIN 72.0 ± 2.48 70.96 ± 1.93 69.76 ± 2.65 70.32 ± 3.27 69.12 ± 2.92 71.04 ± 1.31 70.08 ± 1.21 70.4 ± 2.02 69.12 ± 2.82 70.88 ± 3.98 70.88 ± 2.25 70.8 ± 2.38 IMDB-MUL 49.97 ± 2.16 47.68 ± 4.21 50.13 ± 3.87 50.51 ± 2.92 49.17 ± 4.35 49.76 ± 3.55 47.63 ± 3.45 49.71 ± 2.83 47.79 ± 3.45 49.71 ± 3.7 48.75 ± 3.98 49.49 ± 5.08 REDDIT 76.24 ± 0.54 81.96 ± 1.36 75.68 ± 1.0 74.84 ± 2.68 83.24 ± 1.45 75.56 ± 3.19 82.84 ± 2.54 85.52 ± 1.38 85.12 ± 1.29 74.44 ± 1.74 77.24 ± 1.87 71.28 ± 2.06 ZINC 0.62 ± 0.01 0.57 ± 0.04 0.61 ± 0.01 0.59 ± 0.02 0.51 ± 0.01 0.6 ± 0.01 0.4 ± 0.04 0.34 ± 0.01 0.34 ± 0.02 0.46 ± 0.08 0.36 ± 0.02 0.53 ± 0.04
Table 4: An ablation study compares the performance of CCXN, CWN, CCCN, SCCN, SCCNN, and SCN models on various datasets using two readout strategies, direct readout (DR) and signal down-propagation (SDP). SDP generally enhances CWN performance, whereas the effect of SDP on CCCN, SCCNN, and SCN varies based on their internal signal propagation mechanisms. Means and standard deviations of performance metris are shown. The best results are shown in bold for each model and readout type.
Dataset CCXN CWN CCCN SCCN SCCNN SCN DR SDP DR SDP DR SDP DR SDP DR SDP DR SDP Node-level tasks Cora 86.32 ± 1.22 86.79 ± 1.81 74.95 ± 0.98 86.32 ± 1.38 87.44 ± 1.28 87.68 ± 1.17 80.86 ± 2.16 80.06 ± 1.66 82.19 ± 1.07 80.65 ± 2.39 82.27 ± 1.34 79.91 ± 1.18 Citeseer 72.87 ± 1.13 74.67 ± 2.24 70.49 ± 2.85 75.2 ± 1.82 75.63 ± 1.58 74.91 ± 1.25 69.6 ± 1.83 68.86 ± 2.40 70.23 ± 2.69 69.03 ± 2.01 71.24 ± 1.68 70.4 ± 1.53 Pubmed 88.91 ± 0.47 88.38 ± 0.38 86.94 ± 0.68 88.64 ± 0.36 88.52 ± 0.44 88.67 ± 0.39 88.04 ± 0.51 88.37 ± 0.48 88.18 ± 0.32 87.78 ± 0.58 88.72 ± 0.5 88.62 ± 0.44 Amazon 48.93 ± 0.14 48.34 ± 0.12 45.58 ± 0.25 51.9 ± 0.15 50.55 ± 0.15 50.26 ± 0.17 OOM OOM OOM OOM OOM OOM Empire 80.46 ± 0.23 81.44 ± 0.31 66.13 ± 0.03 81.81 ± 0.62 82.14 ± 0.00 82.51 ± 0.0 88.2 ± 0.22 88.27 ± 0.14 89.15 ± 0.32 88.73 ± 0.12 85.89 ± 0.34 88.79 ± 0.46 Minesweeper $8$ 0.36 89.76 ± 0.32 48.82 ± 0.0 88.62 ± 0.04 89.42 ± 0.00 89.85 ± 0.00 88.85 ± 0.00 89.07 ± 0.25 87.4 ± 0.0 89.0 ± 0.00 90.32 ± 0.11 90.27 ± 0.36 Election 0.39 ± 0.05 0.35 ± 0.02 0.6 ± 0.04 0.34 ± 0.02 0.31 ± 0.02 0.31 ± 0.01 0.53 ± 0.03 0.57 ± 0.02 0.51 ± 0.03 0.56 ± 0.04 0.46 ± 0.04 0.51 ± 0.03 Bachelor 0.33 ± 0.03 0.32 ± 0.03 0.33 ± 0.03 0.33 ± 0.03 0.32 ± 0.02 0.31 ± 0.02 0.36 ± 0.02 0.34 ± 0.02 0.34 ± 0.03 0.34 ± 0.03 0.32 ± 0.02 0.32 ± 0.03 Birth 0.80 ± 0.12 0.74 ± 0.11 0.81 ± 0.11 0.72 ± 0.09 0.71 ± 0.09 0.72 ± 0.05 0.82 ± 0.09 0.83 ± 0.10 0.79 ± 0.12 0.83 ± 0.12 0.71 ± 0.08 0.8 ± 0.11 Death 0.57 ± 0.06 0.54 ± 0.06 0.55 ± 0.05 0.54 ± 0.06 0.54 ± 0.06 0.54 ± 0.06 0.58 ± 0.06 0.56 ± 0.04 0.55 ± 0.05 0.58 ± 0.05 0.52 ± 0.05 0.56 ± 0.05 Income 0.25 ± 0.03 0.25 ± 0.03 0.36 ± 0.04 0.25 ± 0.03 0.23 ± 0.02 0.23 ± 0.02 0.29 ± 0.03 0.29 ± 0.03 0.28 ± 0.03 0.31 ± 0.03 0.25 ± 0.02 0.27 ± 0.02 Migration 0.80 ± 0.11 0.85 ± 0.18 0.9 ± 0.16 0.84 ± 0.13 0.84 ± 0.10 0.84 ± 0.12 0.91 ± 0.18 0.93 ± 0.17 0.90 ± 0.14 0.93 ± 0.17 0.92 ± 0.20 0.96 ± 0.23 Unempl 0.28 ± 0.05 0.27 ± 0.03 0.46 ± 0.04 0.25 ± 0.03 0.24 ± 0.03 0.25 ± 0.03 0.43 ± 0.04 0.47 ± 0.04 0.43 ± 0.04 0.45 ± 0.04 0.38 ± 0.04 0.41 ± 0.03 Graph-level tasks MUTAG 69.79 ± 4.61 74.89 ± 5.51 69.68 ± 8.58 80.43 ± 1.78 80.85 ± 5.42 77.02 ± 9.32 70.64 ± 5.90 73.62 ± 4.41 76.17 ± 6.63 70.64 ± 3.16 71.49 ± 2.43 73.62 ± 6.13 PROTEINS 75.63 ± 2.57 74.91 ± 1.85 76.13 ± 1.80 76.13 ± 2.70 73.55 ± 3.43 73.33 ± 2.30 75.05 ± 2.76 74.34 ± 3.17 74.19 ± 2.86 74.98 ± 1.92 75.27 ± 2.14 74.77 ± 1.69 NCI1 72.43 ± 1.72 74.86 ± 0.82 68.52 ± 0.51 73.93 ± 1.87 76.67 ± 1.48 77.65 ± 1.28 76.42 ± 0.88 76.17 ± 1.39 76.6 ± 1.75 75.6 ± 2.45 75.27 ± 1.57 74.49 ± 1.03 NCI109 73.22 ± 0.48 75.66 ± 1.30 68.19 ± 0.65 73.8 ± 2.06 75.35 ± 1.50 74.83 ± 1.18 75.49 ± 1.39 75.31 ± 1.36 77.12 ± 1.07 75.43 ± 1.94 74.58 ± 1.29 75.7 ± 1.04 IMDB-BIN 70.08 ± 1.21 68.96 ± 2.03 70.4 ± 2.02 69.28 ± 2.57 69.12 ± 2.82 69.44 ± 2.46 70.88 ± 3.98 69.76 ± 3.16 70.88 ± 2.25 69.28 ± 5.69 70.8 ± 2.38 68.64 ± 3.90 IMDB-MUL 47.63 ± 3.45 48.75 ± 3.56 49.71 ± 2.83 49.87 ± 2.33 49.01 ± 2.63 47.79 ± 3.45 49.71 ± 3.70 47.31 ± 3.12 48.75 ± 3.98 46.67 ± 3.13 48.16 ± 2.89 49.49 ± 5.08 REDDIT 74.40 ± 1.50 82.84 ± 2.54 76.20 ± 0.86 85.52 ± 1.38 85.12 ± 1.29 83.32 ± 0.73 74.16 ± 1.54 74.44 ± 1.74 75.56 ± 3.46 77.24 ± 1.87 71.28 ± 2.06 69.68 ± 4.0 ZINC 0.63 ± 0.02 0.40 ± 0.04 0.70 ± 0.0 0.34 ± 0.01 0.35 ± 0.02 0.34 ± 0.02 0.55 ± 0.01 0.46 ± 0.08 0.36 ± 0.01 0.36 ± 0.02 0.59 ± 0.01 0.53 ± 0.04
Table 5: Model sizes corresponding to the best set of hyperparameters
Model GCN GAT GIN AST EDGNN UniGNN2 CWN CCCN CCXN SCN SCCN SCCNN Cora 234.63K 113.61K 105.03K 60.26K 113.29K 109.06K 343.11K 451.85K 735.37K 144.62K 155.88K 164.17K Citeseer 525.06K 558.60K 122.05K 132.87K 258.50K 541.32K 1754.50K 1032.84K 758.41K 737.29K 782.34K 893.13K Pubmed 114.69K 148.23K 53.38K 280.83K 147.59K 114.56K 163.72K 85.76K 277.51K 134.40K 457.99K 605.06K MUTAG 67.97K 22.02K 38.40K 80.77K 5.73K 84.10K 334.72K 284.29K 73.86K 20.03K 398.85K 27.11K PROTEINS 13.19K 10.11K 13.19K 14.34K 5.60K 21.31K 101.12K 34.56K 86.53K 10.24K 397.31K 26.72K NCI1 6.72K 11.20K 154.37K 57.47K 88.19K 104.32K 124.10K 63.87K 15.87K 94.98K 131.84K 188.99K NCI109 23.75K 11.23K 154.50K 221.57K 88.32K 4.61K 412.29K 17.67K 71.36K 26.08K 135.75K 49.54K IMDB-BIN 21.70K 21.83K 9.89K 114.24K 9.86K 100.61K 68.80K 218.24K 19.04K 202.63K 563.07K 285.83K IMDB-MUL 62.08K 6.37K 18.76K 111.30K 27.01K 8.32K 19.68K 45.64K 56.26K 27.94K 545.16K 121.22K REDDIT 13.63K 30.66K 10.08K 106.18K 5.83K 68.10K 26.66K 47.94K 57.54K 7.84K 69.31K 286.59K Amazon 122.37K 156.16K 89.35K 155.91K 122.24K 89.22K 578.95K 310.15K 200.96K OOM OOM OOM Minesweeper 9.28K 5.89K 51.46K 118.02K 21.70K 51.33K 22.24K 35.07K 8.83K 51.97K 25.15K 33.44K Empire 37.39K 41.81K 33.23K 257.17K 41.49K 90.90K 142.74K 86.10K 43.83K 415.89K 612.50K 240.53K Tolokers 84.87K 151.94K 3.75K 217.99K 21.89K 13.57K 12.07K OOM OOM OOM OOM OOM Election 84.22K 151.30K 150.27K 217.34K 21.57K 4.58K 118.08K 234.37K 253.18K 16.64K 11.52K 415.10K Bachelor 17.47K 151.30K 7.81K 316.93K 84.10K 3.55K 43.78K 34.88K 12.86K 27.14K 43.52K 415.10K Birth 4.64K 30.34K 5.70K 30.21K 84.10K 13.25K 26.24K 22.40K 253.18K 103.42K 11.52K 26.98K Death 21.63K 10.18K 7.81K 316.93K 84.10K 51.07K 26.24K 15.58K 12.86K 27.14K 11.52K 415.10K Income 17.47K 151.30K 9.92K 217.34K 21.57K 4.58K 85.18K 12.42K 12.86K 27.14K 268.03K 105.15K Migration 67.71K 10.18K 38.27K 80.64K 21.57K 51.07K 26.24K 15.58K 65.15K 16.64K 11.52K 415.10K Unempl 84.22K 151.30K 117.25K 105.86K 21.57K 5.60K 101.63K 234.37K 286.46K 27.14K 11.52K 105.15K ZINC 22.59K 22.85K 10.40K 106.82K 22.53K 102.14K 88.06K 287.74K 16.48K 24.42K 617.86K 1453.82K Average size 75K ±  [^1] K 89K ± 119K 54K ± 53K 150K ± 88K 59K ± 60K 74K ± 109K 210K ± 367K 170K ± 229K 158K ± 212K 107K ± 172K 263K ± 251K 313K ± 340K
Table 6: TNNs utilized in the experiments and their references
Acronym Neural network name Reference Graph neural networks GAT Graph attention network Veličković et al. [2018] GIN Graph isomorphism network Xu et al. [2019] GCN Semi-Supervised Classification with Graph Convolutional Networks Kipf and Welling [2016] Simplicial complexes SAN Simplicial Attention Neural Networks Giusti et al. [2022] SCCN Efficient Representation Learning for Higher-Order Data with Simplicial Complexes Yang et al. [2022] SCCNN Convolutional Learning on Simplicial Complexes Yang and Isufi [2023] SCN Simplicial Complex Neural Networks Ebli et al. [2020] Cellular complexes CAN Cell Attention Network Giusti et al. [2023b] CCCN Generalized simplicial attention neural networks  [^7] We report the results for the cellular domain of this implementation. Battiloro et al. [2023b] CXN Cell Complex Neural Networks Hajij et al. [2020] CWN Weisfeiler and Lehman Go Cellular: CW Networks Bodnar et al. [2021a] Hypergraphs AllSetTransformer You are AllSet: A Multiset Function Framework for Hypergraph Neural Networks Chien et al. [2021] EDGNN Equivariant Hypergraph Diffusion Neural Operators Wang et al. [2022] UniGNN UniGNN: a Unified Framework for Graph and Hypergraph Neural Networks Huang and Yang [2021]
Appendix D Higher-order domains
A higher-order domain is a generalization of a graph that captures both pairwise and higher-order relationships between entities Hajij et al. [2023]. In the study of topological domains, understanding the distinct properties of various higher-order domains is essential. Two key properties are set-type relations and hierarchical structures, represented by rank functions. Higher-order domains are also called topological domains.
A clarification on terminology. From a mathematical point of view, a higher-order domain is not necessarily a topological space. However, a higher-order domain consists of higher-order cells that capture information about the topology of the underlying domain. From this high-level perspective, higher-order domains (including hypergraphs) are cell structures that encode topological information. In such a context, the terms ‘higher-order domain’ and ‘topological domain’ are used interchangeably. Similarly, the terms ‘higher-order neural network’ and ‘topological neural network’ are used interchangeably.
Definition 4 (Set-type relation)
A relation in a higher-order domain is called a set-type relation if its existence is not implied by another relation in the domain.
Set-type relations emphasize the independence of connections within a domain, allowing for a flexible representation of complex interactions.
Definition 5 (Rank function)
A rank function on a higher-order domain $\mathcal{X}$ is an order-preserving function $rk\colon\mathcal{X}\to\mathbb{Z}_{\geq 0}$; i.e., $x\subseteq y$ implies $rk(x)\leq rk(y)$ for all $x,y\in\mathcal{X}$.
Rank functions introduce a hierarchical organization, facilitating the analysis of nested relationships and enabling pooling operations in deep learning by aggregating information across different levels of the hierarchy.
Over the past few years, four topological domains have found significant applications in TDL: hypergraphs, simplicial complexes, cell complexes, and combinatorial complexes. Note that currently, TopoBenchmarkX library supports simplicial complexes, cell complexes, and hypergraphs, while the rest will be added in future work.
D.1 Hypergraphs
Definition 6 (Hypergraph)
A hypergraph $\mathcal{H}$ on a nonempty set $\mathcal{V}$ is a pair $(\mathcal{V},\mathcal{E}_{\mathcal{H}})$, where $\mathcal{E}_{\mathcal{H}}$ is a non-empty subset of the powerset $\mathcal{P}(\mathcal{V})\setminus\{\emptyset\}$ of the nodes. Elements of $\mathcal{E}_{\mathcal{H}}$ are called hyperedges.
Hypergraphs generalize graphs by allowing edges (called hyperedges) to connect any number of nodes, representing set-type relations without hierarchical structure.
D.2 Abstract simplicial complexes
Definition 7 (Abstract simplicial complex)
An abstract simplicial complex (SC) in a non-empty set $S$ is a pair $SC=(S,\mathcal{X})$, where $\mathcal{X}$ is a subset of $\mathcal{P}(S)\setminus\{\emptyset\}$ such that $x\in SC$ and $y\subseteq x$ imply $y\in SC$. Elements of $\mathcal{X}$ are called simplices.
Simplicial complexes represent hierarchical structures where each simplex (relation) includes all its subsets, but they do not represent set-type relations independently.
D.3 Regular cell complexes
Definition 8 (Regular cell complex)
A regular cell complex is a topological space $S$ with a partition into subspaces (cells) $\{x_{\alpha}\}_{\alpha\in P_{S}}$, where $P_{S}$ is an index set, satisfying the following conditions:
  1. 1.
    $S=\cup_{\alpha\in P_{S}}\text{int}(x_{\alpha})$, where $\text{int}(x)$ denotes the interior of cell $x$.
  2. 2.
    For each $\alpha\in P_{S}$, there exists a homeomorphism $\psi_{\alpha}$, called an attaching map, from $x_{\alpha}$ to $\mathbb{R}^{n_{\alpha}}$ for some $n_{\alpha}\in\mathbb{N}$, called the dimension $n_{\alpha}$ of cell $x_{\alpha}$.
  3. 3.
    For each cell $x_{\alpha}$, the boundary $\partial x_{\alpha}$ is a union of finitely many cells, each having dimension less than the dimension of $x_{\alpha}$.
Cell complexes provide a hierarchical interior-to-boundary structure with clear topological and geometric interpretations, but they are not based on set-type relations.
D.4 Combinatorial complexes
Definition 9 (Combinatorial complex)
A combinatorial complex (CC) is a triple $(\mathcal{V},\mathcal{X},\operatorname{rk})$ consisting of a set $\mathcal{V}$, a subset $\mathcal{X}$ of $\mathcal{P}(\mathcal{V})\setminus\{\emptyset\}$, and a function $\operatorname{rk}\colon\mathcal{X}\to\mathbb{Z}_{\geq 0}$ with the following properties:
  1. 1.
    For all $v\in\mathcal{V}$, $\{v\}\in\mathcal{X}$ and $\operatorname{rk}(v)=0$,
  2. 2.
    The function $\operatorname{rk}$ is order-preserving, which means that if $x,y\in\mathcal{X}$ satisfy $x\subseteq y$, then $\operatorname{rk}(x)\leq\operatorname{rk}(y)$.
Combinatorial complexes combine hierarchical structure with set-type relations, allowing for a flexible and comprehensive representation of higher-order networks.
Appendix E Topological neural networks
Roughly speaking, a topological neural network (TNN) is a neural network that process data on a topological domain. In this section we review the basics of TNNs using the higher-order message passing paradiam introduced in Hajij et al. [2023]. All networks we experiment on in our package fall as a special case of the definition we provide here.
First, we introduce the concept of $k$-cochain spaces, which are essential building blocks for topological neural networks.
Definition 10 (k-cochain spaces)
Let $\mathcal{C}^{k}(\mathcal{X},\mathbb{R}^{d})$ be the $\mathbb{R}$-vector space of functions $\mathbf{H}_{k}\colon\mathcal{X}^{k}\to\mathbb{R}^{d}$ for a rank $k\in\mathbb{Z}_{\geq 0}$ and dimension $d$. $d$ is called the data dimension. $\mathcal{C}^{k}(\mathcal{X},\mathbb{R}^{d})$ is called the $k$-cochain space. Elements $\mathbf{H}_{k}$ in $\mathcal{C}^{k}(\mathcal{X},\mathbb{R}^{d})$ are called $k$-cochains or $k$-signals.
Cochain represent the input topological data (feature vectors) defined on the topological domain. To describe how data is exchanged within a topological domain, we recall the concept of neighborhood functions.
Definition 11 (Neighborhood function)
Let $S$ be a nonempty set. A neighborhood function on $S$ is a function $\mathcal{N}\colon S\to\mathcal{P}(\mathcal{P}(S))$ that assigns to each point $x$ in $S$ a nonempty collection $\mathcal{N}(x)$ of subsets of $S$. The elements of $\mathcal{N}(x)$ are called neighborhoods of $x$ with respect to $\mathcal{N}$.
Definition 12 (TNN)
Let $\mathcal{X}$ be a topological domain. Let $\mathcal{C}^{i_{1}}\times\mathcal{C}^{i_{2}}\times\ldots\times\mathcal{C}^{i_{m}}$ and $\mathcal{C}^{j_{1}}\times\mathcal{C}^{j_{2}}\times\ldots\times\mathcal{C}^{j_{n}}$ be a Cartesian product of $m$ and $n$ cochain spaces defined on $\mathcal{X}$. A topological neural network (TNN) is a function of the form
$TNN\colon\mathcal{C}^{i_{1}}\times\mathcal{C}^{i_{2}}\times\ldots\times\mathcal{C}^{i_{m}}\longrightarrow\mathcal{C}^{j_{1}}\times\mathcal{C}^{j_{2}}\times\ldots\times\mathcal{C}^{j_{n}}.$
A TNN takes as input a vector of cochains $(\mathbf{H}_{i_{1}},\ldots,\mathbf{H}_{i_{m}})$ and produces as output a vector of cochains $(\mathbf{K}_{j_{1}},\ldots,\mathbf{K}_{j_{n}})$. This framework allows for the processing of complex topological data through neural network operations.
Finally, we introduce the concept of higher-order message passing, which is the mechanism by which TNNs can exchange information across different cells in the topological domain:
E.1 Higher-order message passing
Higher-order message passing refers to a computational framework that involves exchanging messages among entities and cells in a higher-order domain using a set of neighborhood functions.
Definition 13 (Higher-order message passing on a topological domain)
Let $\mathcal{X}$ be a topological domain, and let $\mathcal{N}=\{\mathcal{N}_{1},\ldots,\mathcal{N}_{n}\}$ be a set of neighborhood functions defined on $\mathcal{X}$. Consider a cell $x$ and another cell $y\in\mathcal{N}_{k}(x)$ for some $\mathcal{N}_{k}\in\mathcal{N}$. A message $m_{x,y}$ between cells $x$ and $y$ is a computation depending on these two cells or on the data they support. Let $\mathcal{N}(x)$ denote the multi-set $\{\!\!\{\mathcal{N}_{1}(x),\ldots,\mathcal{N}_{n}(x)\}\!\!\}$, and let $\mathbf{h}_{x}^{(l)}$ represent the data supported on the cell $x$ at layer $l$. Higher-order message passing on $\mathcal{X}$, induced by $\mathcal{N}$, is defined through the following update rules:
$m_{x,y}$ $=\alpha_{\mathcal{N}_{k}}(\mathbf{h}_{x}^{(l)},\mathbf{h}_{y}^{(l)}),$ (1)
$m_{x}^{k}$ $=\bigoplus_{y\in\mathcal{N}_{k}(x)}m_{x,y},\quad 1\leq k\leq n,$ (2)
$m_{x}$ $=\bigotimes_{\mathcal{N}_{k}\in\mathcal{N}}m_{x}^{k},$ (3)
$\mathbf{h}_{x}^{(l+1)}$ $=\beta(\mathbf{h}_{x}^{(l)},m_{x}).$ (4)
Here, $\bigoplus$ is a permutation-invariant aggregation function called the intra-neighborhood aggregation of $x$, and $\bigotimes$ is an aggregation function called the inter-neighborhood aggregation of $x$. The functions $\alpha_{\mathcal{N}_{k}}$ and $\beta$ are differentiable functions.
To summarize:
  • $m_{x,y}$ is the message computed from $x$ to $y$ using the function $\alpha_{\mathcal{N}_{k}}$.
  • $m_{x}^{k}$ aggregates all messages from the neighbors $y$ in the neighborhood $\mathcal{N}_{k}(x)$ using the intra-neighborhood function $\bigoplus$.
  • $m_{x}$ further aggregates these results across all neighborhoods $\mathcal{N}_{k}\in\mathcal{N}$ using the inter-neighborhood function $\bigotimes$.
  • $\mathbf{h}_{x}^{(l+1)}$ updates the data on cell $x$ by combining its current data $\mathbf{h}_{x}^{(l)}$ with the aggregated message $m_{x}$ using the function $\beta$.
E.2 Information about neural network models
In this section, we review the architectures of topological neural networks (TNNs) applied to hypergraphs, simplicial complexes, and cellular complexes; for a more comprehensive survey on TNN architectures, refer to Papillon et al. [2023b].
E.2.1 Hypergraphs neural networks
Hypergraph neural networks (HGNNs) have been extensively researched and surveyed. Early methods often reduced hypergraphs to graphs using algorithms such as clique-expansion, which compromised performance by losing structural information Dong et al. [2020]. Current models preserve hypergraph structures and typically employ a two-phase message-passing scheme, transferring information from nodes to hyperedges and then back to nodes.
Standard message passing. Models such as HyperSage, HGNN, Dynamic Hypergraph Neural Network, and UniGCN use standard message passing Wang et al. [2019], Liu et al. [2021], Huang and Yang [2021], Chien et al. [2021]. Notably, some models like UniGCN incorporate fixed weight matrices on top of learnable weights. Other approaches, such as AllSet and EHNN, enhance expressiveness using multi-set functions and sparse symmetric tensors, respectively Chien et al. [2021], Wang et al. [2022].
Attention-based message passing. Models like Dynamic Hypergraph Convolutional Network (SAN), Less is More: Hypergraph Attention Network (SAT), and UniGAT integrate attention mechanisms into the two-phase scheme. Transformer-based variants and multi-head attention are also employed to further enhance performance Wang et al. [2019], Liu et al. [2021].
E.2.2 Simplicial complexes neural networks
Initial research on simplicial complexes focused on signal processing. HodgeNet was one of the first models to learn convolutions on edge features, while subsequent models like SNN and SCCONV generalized convolutions to higher-order features Bodnar et al. [2021a], Bunch et al. [2020], Yang and Isufi [2023].
Standard message passing. Simplicial neural networks (SNNs), SCCONV, and more recent models utilize various neighborhood schemes to enhance message passing flexibility Giusti et al. [2022], Ebli et al. [2020]. High Skip Networks (HSNs) introduce higher-order skip-connections to propagate features through multiple ranks Hajij et al. [2022a].
Attention-based message passing. Models such as SAN, SAT, and SGAT leverage attention mechanisms on simplicial domains, each utilizing unique neighborhood structures and attention coefficients Giusti et al. [2022], Hajij et al. [2022b], Battiloro et al. [2023b].
E.2.3 Cellular complexes neural networks
Building on signal processing advancements, Cell Complex Neural Networks (CXNs) Hajij et al. [2020] were introduced theoretically and later implemented in (CWN) Bodnar et al. [2021a]. The only attention-based cellular network that is Cell Attention Networks (CAN) introduced in Giusti et al. [2023b]. Recently, a cellular transformer architecture (CT) was introduced in Ballester et al. [2024].
Appendix F Descriptive summaries of datasets
Table 7 provides the statistics for each dataset lifted to three topological domains: simplicial complex, cellular complex, and hypergraph. The 00-cell, $1$-cell, $2$-cell, and $3$-cell represents the count of $n$-cell each dataset have after the topology lifting procedure. The clique complex lifting is applied to lift the graph to a simplicial domain with a maximum $n$ equal to three. Cycle-based lifting is employed to lift the graph into the cellular domain (maximum $n$ is equal to two), and $k$-hop lifting is utilized to lift the graph into a hypergraph, where $k=1$.
Table 7: Descriptive summaries of the datasets used in the experiments.
Dataset Domain 00-cell $1$-cell $2$-cell $3$-cell Num. Hyperedges Cora Cellular 2708 5278 2648 0 0 Simplicial 2708 5278 1630 220 0 Hypergraph 2708 0 0 0 2708 Citeseer Cellular 3327 4552 1663 0 0 Simplicial 3327 4552 1167 255 0 Hypergraph 3327 0 0 0 3327 PubMed Cellular 19717 44324 23605 0 0 Simplicial 19717 44324 12520 3275 0 Hypergraph 19717 0 0 0 19717 MUTAG Cellular 3371 3721 538 0 0 Simplicial 3371 3721 0 0 0 Hypergraph 3371 0 0 0 3371 NCI1 Cellular 122747 132753 14885 0 0 Simplicial 122747 132753 186 0 0 Hypergraph 122747 0 0 0 122747 NCI109 Cellular 122494 132604 15042 0 0 Simplicial 122494 132604 183 0 0 Hypergraph 122494 0 0 0 122494 PROTEINS Cellular 43471 81044 38773 0 0 Simplicial 43471 81044 30501 3502 0 Hypergraph 43471 0 0 0 43471 REDDIT-BINARY Cellular 859254 995508 141218 0 0 Simplicial 859254 995508 49670 1303 0 Hypergraph 859254 0 0 0 859254 IMDB-BINARY Cellular 19773 96531 77758 0 0 Simplicial 19773 96531 391991 1694513 0 Hypergraph 19773 0 0 0 19773 IMDB-MULTI Cellular 19502 98903 80901 0 0 Simplicial 19502 98903 458850 2343676 0 Hypergraph 19502 0 0 0 19502 ZINC Cellular 277864 298985 33121 0 0 Simplicial 277864 298985 769 0 0 Hypergraph 277864 0 0 0 277864 Amazon Ratings Cellular 24492 93050 68553 0 0 Simplicial 24492 93050 110765 64195 0 Hypergraph 24492 0 0 0 24492 Minesweeper Cellular 10000 39402 28955 0 0 Simplicial 10000 39402 39204 9801 0 Hypergraph 10000 0 0 0 10000 Roman Empire Cellular 22662 32927 10266 0 0 Simplicial 22662 32927 7168 0 0 Hypergraph 22662 0 0 0 22662 Tolokers Cellular OOM OOM OOM OOM OOM Simplicial OOM OOM OOM OOM OOM Hypergraph 11758 0 0 0 11758 US-county-demos Cellular 3224 9483 6266 0 0 Simplicial 3224 9483 6490 225 0 Hypergraph 3224 0 0 0 3224